OpenStudy (anonymous):

Could someone help quick ? Ill give medal & fan After a discount of 20%, the price of fitness equipment, including tax, was $73.53. The following equation can be used to find x, the price of the equipment without the discount: x - 0.2x + 0.07(x - 0.2x) = 73.53 What is the price of the fitness equipment without the discount? $74.39 $79.06 $85.90 $95.49

OpenStudy (anonymous):

D i think

OpenStudy (anonymous):

yea, D is the closest

OpenStudy (johnweldon1993):

\[\large x - 0.2x + 0.07(x - 0.2x) = 73.53\] Simplify what is in the parenthesis first \[\large x - 0.2x + 0.07(0.8x) = 73.53\] distribute that .07 into the parenthesis... \[\large x - 0.2x + 0.056x = 73.53\] Combine all the like terms on the left \[\large 0.8x + 0.056x = 73.53\] \[\large 0.856x = 73.53\] Divide both sides by \(\large 0.856\) \[\large x = \frac{73.53}{0.856} = 85.90\] Looks like C

OpenStudy (anonymous):

You're awesome lol thanks !

OpenStudy (anonymous):

thats wrong 85.90 x .20 = 17.18 85.90 - 17.8 = 68.1 < discounted price 95.49 x .20 = 19.099 95.49 - 19.099 = 76.391 < discounted price 73.53 is the discounted price the word problem says so we take 76.391 - 73.53 = a difference of 2.861, where as 73.53 - 68.1= 5.43 my answer has less of a difference to the discounted price in the problem then his answer does, so its D

OpenStudy (johnweldon1993):

Good explanation @TheBlackSnow however.... You are forgetting the most important part of the question *including tax* While it IS true that your discounted price is closer to the price they had given...it does not include tax....if it did...your number would be even higher! With my answer...if you were to include tax...you would indeed get closer *or exactly to* the correct discounted price

OpenStudy (anonymous):

let me guess, tax is that 0.02? in the given equation?

OpenStudy (anonymous):

*0.07

OpenStudy (johnweldon1993):

Yeah that looks to be the tax (price with discount + tax on the price with discount = price with discount) yeah .07 is the tax or 7%

OpenStudy (anonymous):

yea, i never use the given equation in a word problem, that's why i guess.

OpenStudy (johnweldon1993):

I don't either...90% of the time it is incorrect...but I figured I would go with this one for the poster :)